What is the smallest positive integer that leaves a remainder of 1 when divided by 12, 25 and 2020?
Answer by ikleyn(52908)

.
What is the smallest positive integer that leaves a remainder of 1 when divided by 12, 25 and 2020?
~~~~~~~~~~~~~~~~~~

Let N be this smallest positive integer number.


Then N-1 is the smallest positive integer number which is a multiple of 12, 25 and 2020.


  12 = 2^2^3;


  25 = 5^2;


  2020 = 20*101 = 2^2*5*101.


Therefore  N-1 = Least Common Multiple of the numbers 12, 25 and 2020 = LCM (12, 25, 2020) = 2^2*3*5^2*101  = 30300.


ANSWER.  The smallest positive integer that leaves a remainder of 1 when divided by 12, 25 and 2020 is 30300+1 = 30301.

Solved.